### v2

@@ III PROBLEM FORMULATION
We extend the single-robot exploration problem formulation in [18] into a multi-robot exploration problem with connectivity constraints. We consider a bounded and unknown environment represented by a 2D occupancy grid map denoted by $\mathcal{M}$. We have $n$ robots $\{1,2,...,n\}$, and each robot maintains its own belief over the exploration map $\mathcal{M}_{i}$. Each map belief consists of unknown region $\mathcal{M}_{u}$ and known region $\mathcal{M}_{k}$, where $\mathcal{M}_{u}\cup\mathcal{M}_{k}=\mathcal{M}$. $\mathcal{M}_{k}$ can be further divided into free area $\mathcal{M}_{f}$ and occupied $\mathcal{M}_{o}$, such that $\mathcal{M}_{f}\cup\mathcal{M}_{o}=\mathcal{M}_{k}$. Each robot is equipped with a 360°$\degree$ lidar, with sensor range $d_{s}$. At the beginning of each exploration mission, we assume that the relative positions of all robots are known.
@@ IV-B Observation Space
The observation for each robot $i$ is $o_{t}=G^{{}^{\prime}t}_{i}$, where $G^{{}^{\prime}t}_{i}=(V^{{}^{\prime}t}_{i},E^{{}^{\prime}t}_{i})$ is the augmented graph modified from $G_{t}$. Each augmented vertex $v^{\prime}_{i,j}=(v_{i,j},u_{i,j},g_{i,j},p_{i,j},s_{i,j})$ comprises five components, of which the first three components belong to the original single-robot formulation [18]. The exploration utility $u_{i,j}$ represents the number of observable frontiers within the node $v_{i,j}$’s line of sight. The guidepost $g_{i,j}$ is a binary value indicating if a node has been visited by robot $i$.
@@
$\small s_{i,j}=\begin{cases}\begin{aligned} &d_{i,j}\frac{\Delta{M_{i,k}}-s_{\text{min}}}{d_{i,k}}+s_{\text{min}}\\ \end{aligned}\vspace{-0.5cm}&\text{if }\Delta{M_{i,k}}\geq\Delta{M_{\text{min}}}\\[20.0pt] 0,&\text{otherwise}\end{cases}$ (3)
@@ IV-D Reward Structure
[…]
@@ V-A1 Encoder
The encoder transforms the explored map into a multi-scale representation through multiple self-attention layers [22]. First, we embed augmented graph vertices $V^{\prime}_{i}$ into d-dimension feature vector $h_{i}^{n}$ As seen in Eq. (4), we obtain the query, key, and value vectors $q_{i}$, $k_{i}$, and $v_{i}$ by multiplying their feature vectors $h_{i}^{q}=h_{i}^{k}=h_{i}^{v}=h_{i}^{n}$ with their learnable weight matrices $W^{Q}$, $W^{K}$, and $W^{V}$ respectively. We then compute the similarity matrix $u_{ij}$ and the attention weights $w_{ij}$ to obtain the enhanced node feature $h^{\prime}_{i}$. Note that each attention layer takes the output of the previous one as input.
@@ VI-B Comparison and Analysis
[…]
@@ VI-D Experimental Validation
We conduct real-world experiments using three four-wheel differential-drive robots in a $\text{25}m\times\text{10}m$ obstacle-rich environment (Fig. 7). Each robot is equipped with an Ouster 32-plane mapping lidar (up to $3.5$m), and with a Doodlelab mesh radio that allows for information hopping using the proximity-based communication model (up to $3.0$m). The robots executes their own decentralized policy using the same trained model used in Sec. V-B2. Throughout the exploration mission, robots repeatedly disconnect and reconnect in an intelligent manner. A snapshot of each robot’s final merged sparse graph representations can be seen in Fig. 8.

### v3

@@ III PROBLEM FORMULATION
[…]
@@ IV-B Observation Space
The observation for each robot $i$ is $o_{t}=G^{{}^{\prime}t}_{i}$, where $G^{{}^{\prime}t}_{i}=(V^{{}^{\prime}t}_{i},E^{{}^{\prime}t}_{i})$ is the augmented graph modified from $G_{t}$. Each augmented vertex $v^{\prime}_{i,j}=(v_{i,j},u_{i,j},g_{i,j},p_{i,j},s_{i,j})$ comprises five components, of which the first three components belong to the original single-robot formulation [18]. The exploration utility $u_{i,j}$ represents the number of observable frontiers within the node $v_{i,j}$’s line of sight. The guidepost $g_{i,j}$ is a binary value indicating if a node has been visited by robot ii.
@@
$\small s_{i,j}=\begin{cases}\begin{aligned} &d_{i,j}\frac{\Delta{M_{i,k}}-s_{\text{min}}}{d_{i,k}}+s_{\text{min}}\\ \end{aligned}\vskip-14.22636pt&\text{if }\Delta{M_{i,k}}\geq\Delta{M_{\text{min}}}\\[20.0pt] 0,&\text{otherwise}\end{cases}$ (3)
@@ IV-D Reward Structure
To achieve better coordination and timely information sharing among multiple robots, we introduce two additional reward components. The first $r_{f}=\Delta(F_{\psi^{t}_{i}},F_{\psi^{t+1}_{i}})$ refers to the increase in the total number of frontier points on the combined map assuming no communication constraints $\mathcal{M}^{t}=(\mathcal{M}_{1}\cup\mathcal{M}_{2}\ldots\mathcal{M}_{n})$. Note that the robots do not directly observe this combined map. Instead, it is only used as privileged information during training to avoid providing incentives for redundant exploration that does not contribute to the team’s overall exploration task. The second $r_{s}=|s_{i,\psi^{t+1}_{i}}|$ refers to the incentive of being on a position along the path of map-surplus utility that robots can observe. This provides dense rewards that encourage robots to pursue other robots for information sharing. The total reward is computed as $\begin{aligned} r_{t}(o_{t},a_{t})=\alpha_{1}\cdot r_{o}+\alpha_{2}\cdot r_{d}+\alpha_{3}\cdot r_{f}+\alpha_{4}\cdot r_{s}+r_{c}.\end{aligned}$
@@ V-A1 Encoder
The encoder transforms the explored map into a multi-scale representation through multiple self-attention layers [22]. First, we embed augmented graph vertices $V^{\prime}_{i}$ into d-dimension feature vector $h_{i}^{n}$ . As seen in Eq. (4), we obtain the query, key, and value vectors $q_{i}$, $k_{i}$, and $v_{i}$ by multiplying their feature vectors $h_{i}^{q}=h_{i}^{k}=h_{i}^{v}=h_{i}^{n}$ with their learnable weight matrices $W^{Q}$, $W^{K}$, and $W^{V}$ respectively. We then compute the similarity matrix $u_{ij}$ and the attention weights $w_{ij}$ to obtain the enhanced node feature $h^{\prime}_{i}$. Note that each attention layer takes the output of the previous one as input.
@@ VI-B Comparison and Analysis
We compare $IR^{2}$ with a pursuit-based approach [13] (Pursuit) and a preplanned-based approach [11] (Preplanned). Pursuit decomposes the map into exploration volumes defined by frontier clusters and allocates them to robots by solving the Vehicle Routing Problem. During exploration, robots decide whether and whom to pursue by weighing the distance cost of deviating from its intended route to obtain map information from others, versus staying on-route to explore the area. On the other hand, Preplanned involves robots agreeing on and adhering to a specified exploration time budget and rendezvous location. For Preplanned, the time budget is pre-set and constant, while the rendezvous position is determined dynamically at a location that minimizes the weighted travel distance for all robots. Note that the original paper [11] includes other robot states intended for task completion, which we leave out for fair comparison. We introduce a 10ss gap between launching each robot to encourage distribution, except for Preplanned which is set at 1ss since it requires all robots to remain in communication range at the beginning.
@@ VI-D Experimental Validation
We conduct real-world experiments using three four-wheel differential-drive robots in a $\text{25}m\times\text{10}m$ obstacle-rich environment (Fig. 7). Each robot is equipped with an Ouster 32-plane mapping lidar (up to 3.53.5m), and with a Doodlelab mesh radio that allows for information hopping using the proximity-based communication model (up to 3.03.0m). The robots executes their own decentralized policy using the same trained model used in Sec. V-B2. Throughout the exploration mission, robots repeatedly disconnect and reconnect in an intelligent manner. A snapshot of each robot’s final merged sparse graph representations can be seen in Fig. 8.